OpenStudy (katie_rogers101):

What is the circumference of the circle if the radius is: 5x/3

OpenStudy (anonymous):

pir^2

OpenStudy (katie_rogers101):

Yes I know the set up I'm just not sure how to solve the problem when it's this way.

OpenStudy (thebossofme191):

oh crap that kinda hard

OpenStudy (unknownrandom):

So you need an x value. Is the x supposed to be pi?

OpenStudy (katie_rogers101):

No I dont think x = pi

OpenStudy (unknownrandom):

Is that the whole question? Are you just trying to create a formula for circumference?

OpenStudy (katie_rogers101):

No I am trying to solve it and That is the whole question.

OpenStudy (unknownrandom):

Well I might be wrong, but I don't think you can solve for the circumference without an x value. You can solve for x, but I think that is about all you can do.

OpenStudy (wolf1728):

Circumference = 2 * PI * radius Circumference = 2 * PI * (5x/3) Circumference = PI * (10x/3)

OpenStudy (wolf1728):

I don't think they want an actual, numerical answer. It's as if they asked "a floor has a width "w" and a length of "l" - What is the area of the floor?" The answer would be (w*l)

OpenStudy (unknownrandom):

@wolf1728 that is what I was thinking.

OpenStudy (wolf1728):

I agree